The degree of ionisation of a 0.1M bromoacetic acid solution is 0.123. Calculate the pH of the solution and the of bromoacetic acid.
1
Expert's answer
2019-09-03T02:33:26-0400

"\space \space \space BrCH_2COOH \to BrCH_2COO^- +H^+"

at t=0 c 0 0

at t=T "c(1-\alpha)" "c \alpha" "c\alpha"

Concentration of "H^+" is given by "c \alpha"

Where "\alpha" is the degree of ionisation.

"[H^+]=c \alpha =0.1 \times 0.123=1.23 \times 10^{-2}" M

"pH=-log[H^+]=-log[1.23 \times 10^{-2}]=1.91"